OpenStudy (anonymous):

Choose the value of the x determinant (Dx) in the following system. 2x + 3y = 5 x – 4y = 7

OpenStudy (anonymous):

det is -11

OpenStudy (lilg132):

you want to solve x? did you understand how to do it using elimination from my previous post?

OpenStudy (anonymous):

btw, det only applies to square matrices, so I'm guessing you mean the left hand side

OpenStudy (anonymous):

yes I understood it a little bit

OpenStudy (mertsj):

||dw:1322351581426:dw

OpenStudy (anonymous):

For a 2x2 matrix \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] the determinant is a*d-b*c

OpenStudy (anonymous):

these are the choices 9 –11 2 –4

OpenStudy (lilg132):

i spent alot of time on that previous post the least you could do is click good answer lol i tried to explain it as clearly as possible which bit did you not understand

OpenStudy (anonymous):

your matrix is \[\left[\begin{matrix}2 & 3 \\ 1 & -4\end{matrix}\right]\]

OpenStudy (anonymous):

lol lilg132..I gave you 2 medals..thanks

OpenStudy (mertsj):

dcrose, that is the matrix of the denominator for both x and y. The x matrix is the same except the x coefficients are replaced by the constants 5 & 7

OpenStudy (anonymous):

im confused...I dont have much time..if someone could possible explain this quickly..

OpenStudy (mertsj):